OpenStudy (anonymous):

A store pays $50.00 for a coat. Its markup is 25%. Later, it puts the coat on sale at 20% off. What is the sale price of the coat?

OpenStudy (tkhunny):

Insufficient Information. There are two kinds of "Markup". Which do we use?

OpenStudy (anonymous):

50 * 1.25 + cost of coat plus 25% markup. $62.50 is cost now. Then store decides to have a sale. figure 205 of the coats cost. .20 * 62.50. This will give you the discount. Now subtract discount from the 62.50 total to get the cost of the coat on sale,

OpenStudy (anonymous):

20% not 205

OpenStudy (tkhunny):

That is one way to do markup. Now the other...

OpenStudy (anonymous):

.20 * 62.50 =12.5 now 62.50 -12.50. you are back to what the store paid for the coat

OpenStudy (anonymous):

What grade are you in? this might make the difference in what tkhunny is thinking.

OpenStudy (tkhunny):

Nope. There are still two ways. Wholesale Markup: $50 + 25% * $50 = Retail Price Retail Markup: $50 = (1-25%)*Retail Price Neither is more difficult. The problem statement just ha to tell you which it is.